OpenStudy (anonymous):

Find all solutions in the interval [0, 2pi). 4sin^(2)x - 4sinx + 1 = 0

OpenStudy (misty1212):

i think this is a perfect square \[(2\sin(x)-1)^2=0\]

OpenStudy (misty1212):

solve \[2\sin(x)-1=0\] do you know how to do it?

OpenStudy (misty1212):

if not, let me know and i can show you

OpenStudy (anonymous):

so from what you have i got x=pi/6, 5pi/6. is that right?

OpenStudy (misty1212):

\[\sin(x)=\frac{1}{2}\] \[x=\frac{\pi}{6}\] and oh yes you are right

OpenStudy (anonymous):

okay i'm just not sure how you got (2sin(x)−1)^2=0 so can you explain how you got there?

OpenStudy (misty1212):

oh ok

OpenStudy (misty1212):

treat \(\sin(x)\) as its own variable like say \(u\) or you can even use \(x\) as long as you switch back, but lets use \(u=\sin(x)\) then you have \[4u^2-4u+1=0\]

OpenStudy (misty1212):

i just recognized this as a perfect square but if you do not, you would try to factor it

OpenStudy (misty1212):

eventually (hope it would be quick) you would see this factors as \((2u-1)(2y-1)\) or \((2u-1)^2\)

OpenStudy (misty1212):

but i knew it because \[a^2-2ab+b^2=(a-b)^2\] and i saw it was of that form

OpenStudy (misty1212):

that is probably more than you wanted to know, but you have a quadratic equation in sine, so you have to solve it like one

OpenStudy (anonymous):

no that was really helpful! i've been stuck on this problem for a little while now lol so thank you so much

OpenStudy (misty1212):

you are welcome, glad to help